# Sparse Autoencoders for Scientifically Rigorous Interpretation of Vision Models

### v1

@@ Abstract
To truly understand vision models, we must not only interpret their learned features but also validate these interpretations through controlled experiments. Current approaches either provide interpretable features without the ability to test their causal influence, or enable model editing without interpretable controls. We present a unified framework using sparse autoencoders (SAEs) that bridges this gap, allowing us to discover human-interpretable visual features and precisely manipulate them to test hypotheses about model behavior. By applying our method to state-of-the-art vision models, we reveal key differences in the semantic abstractions learned by models with different pre-training objectives. We then demonstrate the practical usage of our framework through controlled interventions across multiple vision tasks. We show that SAEs can reliably identify and manipulate interpretable visual features without model re-training, providing a powerful tool for understanding and controlling vision model behavior [^1] Project website: https://osu-nlp-group.github.io/SAE-V.
@@ 1 Introduction
We do not have knowledge of a thing until we have grasped its why, that is to say, its cause.
— Aristotle
Observe Hypothesis Experiment
[…]
Figure 1: We compare the scientific method when applied to genomics and deep learning model interpretation. Grad-CAM (Selvaraju et al., 2017) leverages saliency maps to produce hypothetical explanations for model predictions, but provides no natural way to experimentally validate the hypothesis. In comparison, our proposed use of sparse autoencoders (SAEs) for vision models naturally enables experimental validation via feature suppression. See Section 5 for more examples of experimental validation.
@@
Understanding deep neural networks requires more than passive observation—it demands the ability to test hypotheses through controlled intervention. This mirrors the scientific method itself: true understanding emerges not from mere observation, but from our ability to make and test predictions through controlled experiments. Biologists did not truly understand how genes control traits until they could manipulate DNA and observe the effects. Similarly, understanding vision models requires not just observing behavior, but systematically testing explanations through controlled experiments.
@@
Applying the scientific method to understanding vision models requires three key capabilities. First, we need observable features that correspond to human-interpretable concepts like textures, objects, or abstract properties that appear consistently across images. Biologists need measurable markers of gene expression; we need reliable ways to identify specific visual concepts within our models. Second, we must be able to precisely manipulate these features to test hypotheses about their causal role—like geneticists’ knockout experiments to validate gene function. Finally, methods must work with existing models, just as biological techniques must work with existing organisms rather than requiring genetic redesign.
The scientific method advances understanding through a systematic cycle: observing phenomena, forming hypotheses about their causes, and validating these hypotheses through controlled experiments (Poincaré, 1914; Popper, 1959). This method drove discoveries from genetics to behavioral psychology—Mendel observed patterns in pea plants and tested inheritance theories through careful breeding, while Pavlov’s observation of dogs salivating before feeding led to controlled experiments demonstrating learned associations between stimuli. This sequence of scientific discovery—from observation to hypothesis to experimental validation—provides a proven template that, surprisingly, has not been systematically applied to understanding vision models. Interpretable features enable us to form meaningful hypotheses from observations. Precise control mechanisms let us design experiments to test these hypotheses. Compatibility with existing models allows us to validate our findings on models of interest. The scientific method breaks down if any component is missing: observations become untestable, experiments become uninterpretable, or validation becomes impossible.
Current approaches to understanding vision models fall short of these requirements, preventing rigorous scientific investigation. Feature visualization (Erhan et al., 2009; Mordvintsev et al., 2015a; Olah et al., 2017) methods provide interpretable observations but no mechanism to validate if these observations actually drive model behavior (Geirhos et al., 2024). Adversarial examples (Szegedy, 2013; Goodfellow et al., 2014) enable manipulation but offer no interpretable explanation for how they work. Network dissection (Bau et al., 2017; Hernandez et al., 2021; Kalibhat et al., 2023) attempts to map individual neurons to semantic concepts, but struggles with distributed representations where concepts span multiple neurons or single neurons encode multiple concepts. This makes it difficult to reliably identify and manipulate specific semantic features. This fragmentation—where methods either enable interpretation without control or control without interpretation—undermines our ability to build reliable understanding of these systems
@@
Figure 2: Sparse autoencoders (SAEs) trained on pre-trained ViT activations discover a wide spread of features across both visual patterns and semantic structures. We show eight different features from an SAE trained on ImageNet-1K activations from a CLIP-trained ViT-B/16.
We demonstrate that sparse autoencoders (SAEs; Makhzani & Frey 2013; Subramanian et al. 2018; Huben et al. 2024) enable this kind of rigorous scientific investigation. SAEs transform entangled activation vectors into higher-dimensional sparse representations, in which each nonzero element likely corresponds to a distinct semantic concept. In the case of bird classification, one sparse dimension might indicate blue plumage. Because the SAE is trained to reconstruct the original activations, each element explicitly corresponds to a particular direction in the original dense activation space. This direct mapping allows us to selectively modify that feature—suppressing the blue signal—to observe consequent shifts in model predictions. This unification of interpretation with controlled intervention completes the cycle of observation, hypothesis, and experiment (see Figure 1), and we validate its effectiveness through experiments across multiple core contributions.
@@
First, we show how SAEs enable systematic observation of learned features, revealing fundamental differences between models like CLIP (Radford et al., 2021) and DINOv2 (Oquab et al., 2023) (Section 4). We discover that CLIP learns to recognize country-specific visual patterns, from architectural landmarks to sporting events (Section 4.1) and style-agnostic representations (Section 4.2), suggesting that language supervision leads to rich world knowledge that purely visual training cannot achieve. Just as comparative biology reveals how different environments shape evolution (Grant & Grant, 2006; Losos, 2011; Brawand et al., 2014), our analysis shows how different training objectives lead models to develop qualitatively different internal representations.
@@
Second, we validate our interpretations through controlled experiments across multiple tasks. When models detect features corresponding to bird markings, we confirm their causal role by showing that modifying them predictably changes species classification (Section 5.1). Similarly, with semantic segmentation, we show that identified features enable precise, targeted manipulation: we can suppress specific semantic concepts (like “sand” or “grass”) while preserving all other scene elements, demonstrating both the semantic meaning of our features and their independence from unrelated concepts (Section 5.2). We achieve this through multiple interactive dashboards, enabling readers to readily explore these features. [^2] https://OSU-NLP-Group.github.io/SAE-V#demos
@@
Third, we provide a public, extensible codebase that works with any vision transformer without modification, enabling broad scientific investigation of modern vision models. We demonstrate this flexibility through fine-grained classification and semantic segmentation experiments, with future updates planned.
Through these contributions, we show that rigorous understanding of vision models requires both interpretation and experimental validation, and that SAEs provide a natural concrete implementation of such a framework.
@@ 2 Related Work
Our work connects to three main ideas: methods for interpreting vision models, methods for controlling model behavior, and sparse autoencoder development. We discuss each in relation to our approach.
2.1 Interpretability Methods
Feature visualization methods (Simonyan, 2013; Zeiler, 2014; Mordvintsev et al., 2015b; Olah et al., 2017) reveal interpretable concepts through synthetic images. While these approaches demonstrate that models learn meaningful features, the generated images can be unrealistic or misleading, and the methods cannot validate if these visualizations actually drive model behavior. In contrast, our SAE approach identifies features through real image examples and enables direct testing of their causal influence.
Network dissection (Bau et al., 2017; Zhou et al., 2018; Hernandez et al., 2022; Ghiasi et al., 2022) attempts to map individual neurons to semantic concepts using labeled datasets. However, this approach struggles with distributed representations where concepts span multiple neurons or single neurons encode multiple concepts (polysemanticity). The sparse features learned by our SAEs naturally decompose these distributed representations into interpretable components.
Concept bottleneck models (CBMs; Ghorbani et al. 2019; Koh et al. 2020), prototype-based approaches (Chen et al., 2019; Nauta et al., 2021; Donnelly et al., 2022; Willard et al., 2024), and prompting-based methods (Chowdhury et al., 2025; Paul et al., 2024) explicitly incorporate interpretable concepts into model architecture. While powerful, these methods require model retraining and cannot analyze existing architectures. SAEs can be applied to any pre-trained vision transformer.
@@ 2.1 Interpretability Methods
Follow-up work addresses these shortcomings: Yuksekgonul et al. (2023) convert pre-trained models to CBMS; Schrodi et al. (2024); Tan et al. (2024) develop CBMs with open vocabularies rather than a fixed concept set. In contrast to these works, SAEs reveal a model’s intrinsic knowledge in a task-agnostic manner by decomposing dense activations into sparse, monosemantic features without any retraining. This plug-and-play approach not only faithfully captures the vision model’s internal representations but also enables precise interventions, making SAEs a more flexible tool for validating model interpretation.
Testing with Concept Activation Vectors (TCAV) (Kim et al., 2018) attempts to connect human concepts to model behavior by identifying directions in activation space that correlate with human-specified concepts (like “striped” or “furry”). While TCAV can reveal correlations between activation patterns and semantic concepts, it cannot validate if these concepts actually drive model decisions. This highlights a fundamental limitation of correlation-based interpretation methods: showing that activations correlate with human concepts does not prove those concepts cause model behavior. Our SAE approach moves beyond correlation by enabling controlled experiments—we can actively suppress or enhance specific features and observe the effects on model behavior, providing causal evidence for our interpretations.
2.2 Related Control Methods
Adversarial examples (Szegedy, 2013; Goodfellow et al., 2014; Moosavi-Dezfooli et al., 2016) demonstrate that small, carefully crafted perturbations can dramatically change model predictions. Later work extended this to universal perturbations (Moosavi-Dezfooli et al., 2017) and model reprogramming (Elsayed et al., 2018). However, these perturbations are typically uninterpretable—we cannot understand why they work. Our SAE-based interventions provide interpretable control by manipulating specific semantic features.
Model editing methods for language models (Meng et al., 2022; 2023) enable precise modification of model behavior by directly editing weights or activations. While successful for language tasks, these approaches have not been extended to vision models. Our work provides the first general framework for interpretable editing of vision model behavior.
Recent counterfactual explanation methods (Goyal et al., 2019) attempt to identify minimal input changes that alter predictions. While related, these approaches focus on individual examples rather than discovering and manipulating general features as our method does.
CBMs (Koh et al., 2020) enable editing but only in the last concept layer without spatial resolution. In contrast, our approach allows for localized manipulation, making it widely applicable to various vision tasks.
2.3 Sparse Autoencoders
Makhzani & Frey (2013; 2015) apply $k$-sparse autoencoders to learn improved image representations. Subramanian et al. (2018) apply $k$-sparse autoencoders to static word embeddings (word2vec (Mikolov et al., 2013) and GloVe (Pennington et al., 2014)) to improve interpretability. Zhang et al. (2019); Yun et al. (2021); Bricken et al. (2023); Templeton et al. (2024b); Gao et al. (2024) apply sparse autoencoders to transformer-based language model activations and find highly interpretable features. We extend these ideas to vision, showing that sparsity can reveal interpretable features in visual as well as linguistic domains.
@@ 3 Methodology
3.1 Observations: Model Behaviors
Vision models demonstrate diverse capabilities across multiple tasks. Understanding these systems requires explanations that satisfy three key criteria: they must be human-interpretable, testable through controlled experiments, and enable precise intervention in model behavior.
@@ 3.2 Hypotheses: SAE-Generated Explanations
Figure 3: Given a picture and a set of highlighted patches, we find exemplar images by (1) getting ViT activations for each patch, (2) computing a sparse representation for each highlighted patch (Eqs. 1 and 2), (3) summing over sparse representations, (4) choosing the top $k$ features by activation magnitude and (5) finding existing images that maximize these features.
Sparse autoencoders generate testable hypotheses by decomposing dense activation vectors into sparse feature vectors. Given an $d$-dimensional activation vector $\mathbf{x}\in\mathbb{R}^{d}$ from an intermediate layer $l$ of a vision transformer, an SAE maps $\mathbf{x}$ to a sparse representation $f(\mathbf{x})$ (Eqs. 1 and 2) and reconstructs the original input (Eq. 3). We use ReLU autoencoders (Bricken et al., 2023; Templeton et al., 2024b):
@@
$\mathcal{L}(\theta)=||\mathbf{x}-\mathbf{\hat{x}}||_{2}^{2}+\lambda\mathcal{S}(f(\mathbf{x}))$ (4)
@@
We train on $N$ randomly sampled patch activation vectors from the residual stream of layer $l$ in a vision transformer. Following prior work (Templeton et al., 2024b), we subtract the mean activation vector and normalize activation vectors to unit norm before training. Detailed reproduction instructions are available. Given a pre-trained ViT like CLIP or DINOv2, an image, and one or more patches of interest, we leverage a trained SAE to find similar examples (Figure 3).
@@ 3.3 Experiments: Testing Through Control
    Calculate reconstruction error (Templeton et al., 2024b): $\mathbf{e}=\mathbf{x}-\mathbf{\hat{x}}$.
@@
In plain language: We start by converting an image into a set of activation vectors—one per patch—that capture the ViT’s internal representation. Then, we encode these vectors into a sparse representation that highlights the key features, while also tracking the small differences (errors) between the sparse form and the original. Next, we deliberately tweak the sparse representation to modify a feature of interest. After reconstructing the modified activation vectors (and adding back the previously captured details), we pass them through the task-specific decoder (e.g., for classification, segmentation, or another vision task) to see how the output changes. By comparing the original and altered outputs, we can determine whether and how the targeted feature influences the model’s behavior.

### v2

# Interpretable and Testable Vision Features via Sparse Autoencoders

@@ Abstract
To truly understand vision models, we must not only interpret their learned features but also validate these interpretations through controlled experiments. While earlier work offers either rich semantics or direct control, few post-hoc tools supply both in a single, model-agnostic procedure. We use sparse autoencoders (SAEs) to bridge this gap; each sparse feature comes with real-image exemplars that reveal its meaning and a decoding vector that can be manipulated to probe its influence on downstream task behavior. By applying our method to widely-used pre-trained vision models, we reveal meaningful differences in the semantic abstractions learned by different pre-training objectives. We then show that a single SAE trained on frozen ViT activations supports patch-level causal edits across tasks (classification and segmentation) all without retraining the ViT or task heads. These qualitative, falsifiable demonstrations position SAEs as a practical bridge between concept discovery and causal probing of vision models. We provide code, demos and models on our project website: https://osu-nlp-group.github.io/saev.
@@ 1 Introduction
Understanding deep neural networks requires more than passive observation; it demands the ability to test hypotheses through controlled intervention. This mirrors the scientific method itself: true understanding emerges not from mere observation, but from our ability to make and test predictions through controlled experiments. Biologists did not truly understand how genes control traits until they could manipulate DNA and observe the effects. Similarly, understanding neural networks requires not just observation, but systematically testing explanations through controlled experiments.
@@
Applying the scientific method to understanding vision models requires three key capabilities. First, we need observable features that correspond to human-interpretable concepts like textures, objects, or abstract properties. Biologists need measurable markers of gene expression; we need reliable ways to identify specific visual concepts within our models. Second, we must be able to precisely manipulate these features to test hypotheses about their causal role, like geneticists’ knockout experiments to validate gene function. Finally, methods must work with existing models, just as biological techniques must work with existing organisms rather than requiring genetic redesign.
The scientific method of observe, hypothesize and intervene maps neatly onto machine learning interpretability: observe model behavior, hypothesize why a model makes a particular prediction, then perturb the system to see how behavior changes [poincare1914science, popper1959logic]. Saliency and attribution maps [simonyan2013deep, selvaraju2017gradcam], feature-visualization [mordvintsev2015deepdream, olah2017feature] and network-dissection [bau2017broden] methods link model activations to human concepts, acting as a hypothesized explanation for model behavior. Yet translating those hypotheses into controlled tests is fraught with challenges: masking-based ablations risk distribution shift [hooker2019benchmark]; other surveys conclude that quantitative faithfulness tests remain elusive [adebayo2018sanity, distill2020attribution]. Concept activation vectors [kim2018tcav, goyal2019counterfactual] translate human concepts into testable relationships, but require hand-labeled concept sets. Concept-bottleneck models embed editable concepts directly in the architecture [koh2020concept], but doing so demands retraining the model. Consequently, few post-hoc techniques offer (i) human-interpretable features, (ii) a natural intervention on those features, and (iii) compatibility with pre-trained vision models.
We show in this work that sparse autoencoders (SAEs) offer a natural solution to this problem. SAEs transform dense, entangled activation vectors into higher-dimensional sparse representations in which each element likely corresponds to a distinct semantic concept. Because the SAE is trained to reconstruct the original activations, each element also explicitly corresponds to a particular direction in the original dense activation space. Thus, each element in the sparse representation has both (1) exemplar patches, yielding a semantic visual description, and (2) a corresponding decoding vector in the original space, yielding a precise method for causal interventions. For example, a blue-jay wing activates an element whose exemplar patches are nothing but blue feathers. Suppressing model activations in that direction pushes the classifier away from blue-feathered species, completing the intervene step and confirming the feature’s causal role (see Fig.˜6 for an example). In this work, we demonstrate that SAEs are a natural, intuitive post-hoc method for the complete observe–hypothesize–intervene cycle.
@@
Figure 1: Sparse autoencoders (SAEs) trained on pre-trained ViT activations discover a wide spread of features across both visual patterns and semantic structures. We show eight different features from an SAE trained on ImageNet-1K activations from a CLIP-trained ViT-B/16. Colored patches mark where the SAE features fire within an image; each SAE feature fires on on semantically consistent but visually diverse patches.
@@
First, we use SAEs to survey learned features in CLIP [radford2021learning] and DINOv2 [oquab2023dinov2] and illustrate consistent differences (Section˜4). We discover that CLIP learns to recognize country-specific visual patterns (Section˜4.1) and style-agnostic representations (Section˜4.2), suggesting that language supervision leads to rich world knowledge that purely visual training cannot achieve. Just as comparative biology reveals how different environments shape evolution [grant2006evolution, losos2011lizards, brawand2014genomic], our analysis shows that different training objectives lead to qualitatively different internal representations.
@@
Second, we validate our SAE interpretations on two canonical tasks. When models detect features corresponding to bird markings, we confirm their causal role by showing that modifying them predictably changes species classification (Section˜5.1). Similarly, with semantic segmentation, we show that identified features enable precise, targeted manipulation: we can suppress specific semantic concepts (like “sand” or “grass”) while preserving all other scene elements, demonstrating both the semantic meaning of our features and their independence from unrelated concepts (Section˜5.2). Interpretability is still, as [olah2024qualitative] put it, “an early, messy, un-established science in which low-hanging qualitative structure is often more trustworthy than synthetic summary scores” [olah2024qualitative]. We therefore follow their advice, and the cautions of [adebayo2018sanity] and [hooker2019benchmark] by prioritizing human-falsifiable interventions over aggregate metrics for the present work. Our public, interactive webapps enable readers to explore these features and perform their own causal edits, turning our claims into falsifiable, interactive evidence. [^1] https://anonymous-saev.github.io/saev/demos/semseg and https://anonymous-saev.github.io/saev/demos/classification
@@
Third, we provide a public, extensible codebase that works with any vision transformer, enabling broad investigation of modern vision models. We demonstrate this flexibility through fine-grained classification and semantic segmentation experiments, with future updates planned.
Through these contributions, we demonstrate that sparse autoencoders are a practical method to link semantic feature discovery and interactive causal testing in modern vision models.
@@ 2 Related Work
Saliency methods and attribute maps [zhou2016learning, sundararajan2017axiomatic, selvaraju2017gradcam, wang2020scorecam] highlight relevant parts of images for a given prediction; these methods excel at observation but require masking to test causality, which can introduce distribution shift [adebayo2018sanity, hooker2019benchmark]. Feature visualization methods [simonyan2013deep, zeiler2014visualizing, mordvintsev2015inceptionism, olah2017feature] reveal interpretable concepts through synthetic images. While these approaches demonstrate that models learn meaningful features, the generated images are unrealistic and we cannot validate if these visualizations actually drive model behavior. In contrast, SAEs identify features through real image examples and enables direct testing of their causal influence. Network dissection [bau2017broden, zhou2018interpreting, hernandez2021natural, ghiasi2022vision] attempts to map individual neurons to semantic concepts using labeled datasets. However, this approach struggles when single neurons encode multiple concepts [goh2021multimodal]. In contrast, SAEs (and dictionary learning in general) are explicitly designed to naturally decompose polysemantic representations into single-concept interpretable components. Concept activation vectors [[, CAVs,]]kim2018tcav,goyal2019counterfactual,ghorbani2019towards,singla2021causal provide a natural intervention once a concept dataset exists, but they depend on hand-labeled positive examples and deliver one vector per concept, limiting exploratory use. We train SAEs on pre-trained models and explore the models’ visual vocabularies without labeling any examples.
@@
Concept bottleneck models (CBMs; [ghorbani2019towards, koh2020concept]), prototype-based approaches [chen2019looks, nauta2021looks, donnelly2022deformable, willard2024looks], and prompting-based methods [paul2024simple, chowdhury2025prompt] explicitly incorporate interpretable concepts into model architecture. While powerful, these methods require specific pre-training model architectures and objectives and cannot analyze existing pre-trained models. SAEs can be applied to and analyze any pre-trained vision transformer. Follow-up work addresses these shortcomings: [yuksekgonul2023posthoc] convert pre-trained models to CBMS; [schrodi2024concept, tan2024explain] develop CBMs with open vocabularies rather than a fixed concept set. In contrast to these works, SAEs reveal a model’s intrinsic knowledge in a task‐agnostic manner by decomposing dense activations into sparse, monosemantic features without any retraining. This plug‐and‐play approach not only faithfully captures the vision model’s internal representations but also enables precise interventions, making SAEs a more flexible tool for validating model interpretation.
Most similar to our approach is DN-CBM [rao2024discover], which trains sparse autoencoders on CLIP embeddings and then names units via text similarity before training a new linear concept bottleneck model. Our work differs in three respects: (i) we leave the backbone and task head frozen, editing activations directly; (ii) we validate edits on both classification and semantic segmentation, demonstrating task agnosticism; (iii) we analyze two pre-trained models (CLIP vs DINOv2) to study representation differences.
Sparse autoencoders are a popular dictionary learning method for extracting concept libraries. [makhzani2013k, makhzani2015winner] apply $k$-sparse autoencoders to learn improved image representations. [subramanian2018spine] apply $k$-sparse autoencoders to static word embeddings (word2vec [mikolov2013word2vec] and GloVe [pennington2014glove]) to improve interpretability. [zhang2019word, yun2021transformer, bricken2023monosemanticity, templeton2024scaling, gao2024scaling] apply sparse autoencoders to transformer-based language model activations and find highly interpretable features. [lim2025sparse] apply SAEs to vision models to explain prompt-based tuning methods and [thasarathan2025universal] apply SAEs to vision models for cross-model concept alignment. We aim to convincingly demonstrate that feature descriptions (exemplar images) aligned with causal interventions (decoder vectors) are a natural fit for the observe-hypothesize-intervene loop.
@@ 3 Methodology
Vision models demonstrate remarkable capabilities, but understanding their behavior requires rigorous scientific investigation. We present a framework that enables systematic observation, hypothesis formation, and experimental validation using sparse autoencoders (SAEs).
3.1 Observations: Feature Discovery
Given a pre-trained vision model, we first need reliable ways to observe its internal representations. Traditional approaches like visualizing individual neurons or studying attention patterns provide limited insight. Instead, we train SAEs on activation vectors from intermediate model layers, enabling systematic observation of learned features.
Figure 2: Given a picture and a set of highlighted patches, we find exemplar images by (1) getting ViT activations for each patch, (2) computing a sparse representation for each highlighted patch (Eqs.˜1 and 2), (3) summing over sparse representations, (4) choosing the top $k$ features by activation magnitude and (5) finding existing images that maximize these features.
@@ 3.2 Hypotheses: SAE-Generated Explanations
Sparse autoencoders generate testable hypotheses by decomposing dense activation vectors into sparse feature vectors. Given an $d$-dimensional activation vector $\mathbf{x}\in\mathbb{R}^{d}$ from an intermediate layer $l$ of a vision transformer, an SAE maps $\mathbf{x}$ to a sparse representation $f(\mathbf{x})$ (Eqs.˜1 and 2) and reconstructs the original input (Eq.˜3). We use ReLU autoencoders [bricken2023monosemanticity, templeton2024scaling]:
@@
[…]
@@
We train on $N$ randomly sampled patch activation vectors from the residual stream of layer $l$ in a vision transformer. Following prior work [templeton2024scaling], we subtract the mean activation vector and normalize activation vectors to unit norm before training. See Section˜8 for additional details. Given a pre-trained ViT like CLIP or DINOv2, an image, and one or more patches of interest, we leverage a trained SAE to find similar examples (Fig.˜2).
@@ 3.3 Experiments: Testing Through Control
    Calculate reconstruction error [templeton2024scaling]: $\mathbf{e}=\mathbf{x}-\mathbf{\hat{x}}$.
@@
In plain language: We start by converting an image into a set of activation vectors (one per patch) that capture the ViT’s internal representation. Then, we encode these vectors into a sparse representation that highlights the key features, while also tracking the small differences (errors) between the sparse form and the original. Next, we deliberately tweak the sparse representation to modify a feature of interest. After reconstructing the modified activation vectors (and adding back the previously captured details), we pass them through the task-specific decoder (e.g., for classification, segmentation, or another vision task) to see how the output changes. By comparing the original and altered outputs, we can determine whether and how the targeted feature influences the model’s behavior.